### v1

@@ 5.3.2 Experimental results
Data collection is done at a control and sampling frequency of 200 Hz, which is higher than in the simulation environment, but necessary for the agile maneuvering of the real quadrotor. The size of the datasets are reported in Table 10. For data collection and to conduct the experiments, we use the same reference paths as in the simulation environment, executed by the Mellinger controller [29] for reference tracking.
@@
Train 123 197123197123\,197123 197 228 323228323228\,323228 323
Validation 21 7402174021\,74021 740 40 2924029240\,29240 292
Test $480$ 3 00030003\,0003 000
@@ 6 Conclusion
To show that (25) is an exact embedding of (21), we employ function factorization through the second fundamental theorem of calculus (FTC), extending the approach in [12] to systems with process noise. Based on (22), we have the following decomposition:
(55) $x_{k+1}=f_{\mathrm{d}}(x_{k},u_{k},e_{k})=f(x_{k})+g(x_{k},u_{k})+d(x_{k},u_{k},e_{k})\vspace{-.2cm}$
@@
(56) $\Phi(x_{k+1})=\Phi(f(x_{k}))=A\Phi(x_{k}).\vspace{-.15cm}$
Step 2: Embedding the control input part: Take $e_{k}=0$, implying $d(x_{k},u_{k},0)=0$ such that $x_{k+1}=f(x_{k})+g(x_{k},u_{k})$. Using the results in [12] and (56), the exact lifted form of $x_{k+1}=f(x_{k})+g(x_{k},u_{k})$ is:
[…]
@@
we can apply the proof of Theorem 2 in [12]: in Eq. (43) in [12], chose $q_{k+1}=x_{k+1}$ (which is expanded as (58)) and $p_{k+1}=f(x_{k})+g(x_{k},u_{k})$, giving an exact lifted form that includes the effect of noise as:
[…]
By applying the exact factorization Lemma 1 in [12], we get:
@@
[…]
@@
with $B(z_{k},u_{k}):=B_{\mathrm{x}}(\Phi^{\dagger}(z_{k}),u_{k})$, $K(z_{k},u_{k},e_{k}):=K_{\mathrm{x}}(\Phi^{\dagger}(z_{k}),u_{k},e_{k})$.

### v3

@@ 5.3.2 Experimental results
Data collection is done at a control and sampling frequency of 200 Hz, which is higher than in the simulation environment, but necessary for the agile maneuvering of the real quadrotor. The size of the datasets are reported in Table 10. For data collection and to conduct the experiments, we use the same reference paths as in the simulation environment, executed by the Mellinger controller [mellinger_control] for reference tracking.
@@
Train 123 197123\,197 228 323228\,323
Validation 21 74021\,740 40 29240\,292
Test 480480 3 0003\,000
@@ 6 Conclusion
To show that (25) is an exact embedding of (21), we employ function factorization through the second fundamental theorem of calculus (FTC), extending the approach in [Aut_Iacob_inputs] to systems with process noise. Based on (22), we have the following decomposition:
(55) $x_{k+1}=f_{\mathrm{d}}(x_{k},u_{k},e_{k})=f(x_{k})+g(x_{k},u_{k})+d(x_{k},u_{k},e_{k})\vskip-5.69046pt$
@@
(56) $\Phi(x_{k+1})=\Phi(f(x_{k}))=A\Phi(x_{k}).\vskip-4.26773pt$
Step 2: Embedding the control input part: Take $e_{k}=0$, implying $d(x_{k},u_{k},0)=0$ such that $x_{k+1}=f(x_{k})+g(x_{k},u_{k})$. Using the results in [Aut_Iacob_inputs] and (56), the exact lifted form of $x_{k+1}=f(x_{k})+g(x_{k},u_{k})$ is:
(57) $\Phi(x_{k+1})-\underbrace{\Phi(f(x_{k}))}_{A\Phi(x_{k})}=\underbrace{\left(\int^{1}_{0}\frac{\partial\Phi}{\partial x}(f(x_{k})+\lambda g(x_{k},u_{k}))\operatorname{d\!}\lambda\right)g(x_{k},u_{k})}_{\tilde{B}_{\mathrm{x}}(x_{k},u_{k})}.$
@@
we can apply the proof of Theorem 2 in [Aut_Iacob_inputs]: in Eq. (43) in [Aut_Iacob_inputs], chose $q_{k+1}=x_{k+1}$ (which is expanded as (58)) and $p_{k+1}=f(x_{k})+g(x_{k},u_{k})$, giving an exact lifted form that includes the effect of noise as:
(59) $\Phi(x_{k+1})-\underbrace{\Phi(f(x_{k})+g(x_{k},u_{k}))}_{A\Phi(x_{k})+\tilde{B}_{\mathrm{x}}(x_{k},u_{k})}=\\ =\underbrace{\left(\int^{1}_{0}\frac{\partial\Phi}{\partial x}(f(x_{k})+g(x_{k},u_{k})+\lambda d(x_{k},u_{k},e_{k}))\operatorname{d\!}\lambda\right)d(x_{k},u_{k},e_{k})}_{\tilde{K}_{\mathrm{x}}(x_{k},u_{k},e_{k})}.$
By applying the exact factorization Lemma 1 in [Aut_Iacob_inputs], we get:
@@
$B_{\mathrm{x}}(x_{k},u_{k})=\int^{1}_{0}\frac{\partial\tilde{B}_{\mathrm{x}}}{\partial u}(x_{k},\lambda u_{k})\operatorname{d\!}\lambda,\;\;K_{\mathrm{x}}(x_{k},u_{k},e_{k})=\int^{1}_{0}\frac{\partial\tilde{K}_{\mathrm{x}}}{\partial e}(x_{k},u_{k},\lambda e_{k})\operatorname{d\!}\lambda.$
@@
with $B(z_{k},u_{k})\mathrel{\mathop{\ordinarycolon}}=B_{\mathrm{x}}(\Phi^{\dagger}(z_{k}),u_{k})$, $K(z_{k},u_{k},e_{k})\mathrel{\mathop{\ordinarycolon}}=K_{\mathrm{x}}(\Phi^{\dagger}(z_{k}),u_{k},e_{k})$.